OpenStudy (anonymous):

NEED HELP FAST AND WILL GIVE MEDAL :C CAN YOU HELP ME FIND THE SIMPLIFIED FORM OF THESE PROBLEMS! 1. (x^-4)^5 x^20 1/x^20 x/20 1/x^1,024 2. (-x^6)^3 -x^18 -x^9 x^3 x^18 3. (q^9)^2 2q^18 q^18 q^81 q^11 4. (x^-3)^-5x^6 x^21 x^-21 x^48 x^-48 5. (5t^3)^-4 625/t^12 20/t7 1/625t^12 1/20t^7 6. (x/2y^5)^-2 x^2/4y^10 4y^25/x^2 4y^10/x^2 x^2/4y^25 7. (x^3/2y^6)^-3 x^9/8y^18 8y^18/x^9 6y^9/x^6 x^6/-6y^9 I also need help with this: 8. What is the value of 12x^-3y^-1 for x = -1 and y = 5? -12/5 -180 7,500 -5/12

OpenStudy (anonymous):

PLEASE HELP ME

OpenStudy (noorik.):

1) 1/x^20 2)-x^18 3) q^18 4)x^21 5)625/t^12 6)4y^10/x^2 7)8y^18/x^9

OpenStudy (anonymous):

Thank you so much! What about 8? :D

OpenStudy (noorik.):

12(-1)^-3 (5)^-1 = -12* 1/5 = -12/5 i think it's like that

OpenStudy (anonymous):

Thanks so much :D

OpenStudy (noorik.):

you r welcome